OpenStudy (anonymous):

What is the resistance of a 75W light bulb connected to its proper source voltage of 110V?

OpenStudy (irishboy123):

clues: 1) 75W light bulb [ie Power] 2) proper source voltage of 110V [ie voltage] from Ohm's Law, we have a lot to play with: \(V = I R; \ I = V/R \) and, if we remember a fundamental idea: \(P = I \times V\), then we can manipulate that using Ohm's Law to get: \(P = V^2 / R\) so we have an equation that allows us to relate resistance to voltage and power. and we can begin to understand why that relationship exists.